# EMMOE: A Comprehensive Benchmark for Embodied Mobile Manipulation in Open Environments

### v1

@@ A.3 LLMs for Task Planning
Typical usages of LLM for task planing include treating LLM as a translator or a planner. There are also some studies combining LLMs with traditional PDDL (Guan et al., 2023; Valmeekam et al., 2024; Silver et al., 2024; Zhou et al., 2024), in which LLM will be treated as a translator between the real-world and specific domain planner. But this method is limited by the performance of the domain planner and can’t leverage the powerful commonsense reasoning capabilities of LLMs to assist in planning. When LLM is treated as a planner, discrepancies between LLM’s outputs and real-world conditions always lead to execution failures. LLM-Planner (Song et al., 2023) make a straightforward re-plan. Self-Refine (Madaan et al., 2024) use single LLM as generator and evaluator. Reflexion (Shinn et al., 2024) treats LLM as the Actor and the evaluator as the Critic. ViLA (Lin et al., 2024) utilizes GPT-4V (Yang et al., 2023b) to obtain visual feedback. However, self-improvement relies heavily on prompt design and intrinsic capabilities of LLMs. If errors unrelated to planning occur, LLMs may struggle to self-correct. Inner Monologue (Huang et al., 2022) and RoCo (Mandi et al., 2024) utilizes external collision detection and feedback. DoReMi (Xie et al., 2024) sets pre-defined constrains. Nevertheless, LLMs may make same mistakes in similar situations as the model weights are not changed. SayCan (Ahn et al., 2022) trains a value function to consider both generated actions and their values. Remember (Zhang et al., 2024a) builds a memory module and retrieves similar state-action pairs. Retroformer (Yao et al., 2023) learns a retrospective model via policy gradient optimization. While RL-based adaptation mechanisms are able to adjust actions before execution, defining and training an effective value function or reward model is highly challenging. The recently popular DPO (Rafailov et al., 2024) algorithm greatly simplifies this process by requiring only a preference dataset. In our  HomieBot, we use DPO for model alignment, CoT (Wei et al., 2022) and self-reflection for decision-making. Additionally, error detection and feedback mechanisms are applied during low-level execution.
@@ B.1 Data Collection
We first randomly sample episode information provided by Replica Challenge (Szot et al., 2021) to build the task scenario, then we use the Fetch robot to collect EMMOE-100 in Habitat-lab v0.2.3. To facilitate data collection, we modify the original interaction script, and new interface can be seen in Fig. B1. The interface provides both third-person and first-person view observation to facilitate data collection, third-person observation is used to assist with collection, only first-person observation with 256*256 resolution will be saved. Notably, we only use the scene information to collect environment data, other functions and metrics in Replica Challenge are irrelevant to our work.
@@ E.3 Models
RT-1-X(Brohan et al., 2022) RGB & Instructions Manipulation Picking & Placing
Octo(Team et al., 2024b) RGB & Instructions Manipulation Opening & Closing
NoMaD(Sridhar et al., 2024) RGB & Goal-Image Image-Navigation Navigate to Spot & Large Object
PixNav(Cai et al., 2024) RGB & Goal-Name Pixel-Navigation Navigate to Object
@@
(Gu et al., 2022) can flexible interact with target objects from various locations based on the integration of manipulative skills and mobility, while navigational skills are designed to accommodate multiple endpoints, ultimately leading to successful operations. Specifically, M3 implements these concepts by emphasizing mobile manipulation skills over fixed skills and training navigational skills using area targets rather than point targets.
@@
(Padalkar et al., 2023)) architecture utilizes image and text instructions as inputs, and generates discrete end-effector actions as outputs. Specifically, RT-1-X is a transformer-based model that guides robotic arms to complete various manipulation tasks. RT-1-X is an extension of the RT-1 ( (Brohan et al., 2022)) model, which is designed for robot control and trained on a large-scale robot dataset.
@@
(Team et al., 2024b)) is an open-source, general-purpose policy for robotic manipulation based on transformers. It supports flexible task and observation definition and can be quickly integrated into new observation and action spaces.
@@
(Sridhar et al., 2024)) trains a single diffusion strategy for goal-oriented navigation and goal-independent exploration, the first one is to reach user-specified goals after localization and the second one is to search new environments. The method is instantiated using a transformer-based large-scale policy trained on data from various ground robots.
@@
(Cai et al., 2024)) is a pixel-guided navigational skill. It designs an LLM-based planner that utilizes common sense between objects and rooms to select the optimal waypoints, which are then executed by a pixel navigation strategy to achieve long-line-of-sight navigation. In this pipeline, we use its ability of finding the optimal waypoint and pixel navigation to navigate to some specific small object such as lemon and sugar box.
@@ F.1 SFT Augmentation
To expand the original dataset size, we first use GPT-4o (Achiam et al., 2023) to regenerate text descriptions. Here is the regeneration code clip, we just show how to regenerate task descriptions, but the regeneration of subtask analysis uses the same template.
@@ G.1 Training Parameters
We use Video-LLaVA-7B (Zhang et al., 2023) as our base model, we also use the training scripts they provide and partial parameters for $sft$ are as follows.
@@ H.1 Baseline Setup
Despite lowering the output format standards, the output from 7B-sized models still fails to meet our least requirements. They either do not output single-step subtasks or the subtask format is far from requirements. This issue is difficult to resolve by merely adjusting prompts. Therefore, we leverage the in-context learning abilities of these models by providing an output template example before each inference. Here, we provide the inference template for Qwen2-VL (Wang et al., 2024b) MiniCPM-V 2.6 (Yao et al., 2024) respectively.
@@ H.2 Results
GPT-4o(Achiam et al., 2023) 3.97 0.79 0.79 0 5.56 44.44 0 44.44 1.59 17.46 19.05 15.87 15.08 30.95 30.29
Gemini-1.5-Pro(Team et al., 2024a) 3.85 3.85 0 7.69 15.38 48.08 0 48.08 0 17.31 17.31 15.38 3.85 19.23 21.80
Qwen2-VL-7B(Wang et al., 2024b) 0 0 0 0 0 100 0 0 0 0 0 0 0 0 20
MiniCPM-V 2.6(Yao et al., 2024) 0 0 0 0 0 100 0 0 0 0 0 0 0 0 6.67
@@
GPT-4o(Achiam et al., 2023) 6.87 0.12 0.69 3.65 11.34 8.41 0.06 8.47 0.57 64.88 65.45 13.99 0.75 14.74 73.61
Gemini-1.5-Pro(Team et al., 2024a) 7.48 1.52 2.41 6.45 17.86 9.41 0 9.41 0 47.86 47.86 22.76 2.10 24.86 68.38
Qwen2-VL-7B(Wang et al., 2024b) 2.17 9.49 0.99 3.56 16.21 7.71 0 7.71 4.74 54.35 59.09 16.40 0.59 17.00 27.74
MiniCPM-V 2.6(Yao et al., 2024) 8.58 0.80 0.92 1.72 12.01 7.78 0 7.78 3.49 65.39 68.88 10.87 0.46 11.33 31.08
@@
GPT-4o(Achiam et al., 2023) 5 1 1 0 7/126 56 0 56/126 2 22 24/126 20 19 39/126 126/416
Gemini-1.5-Pro(Team et al., 2024a) 4 4 0 8 16/104 50 0 50/104 0 18 18/104 16 4 20/104 104/477
Qwen2-VL-7B(Wang et al., 2024b) 0 0 0 0 0/9 9 0 9/9 0 0 0/9 0 0 0/9 9/45
MiniCPM-V 2.6(Yao et al., 2024) 0 0 0 0 0/1 1 0 0/1 0 0 0/1 0 0 0/1 1/15
@@
GPT-4o(Achiam et al., 2023) 228 4 23 121 376/3317 279 2 281/3317 19 2152 2171/3317 464 25 489/3317 3317/4506
Gemini-1.5-Pro(Team et al., 2024a) 217 44 70 187 518/2900 273 0 273/2900 0 1388 1388/2900 660 61 721/2900 2900/4241
Qwen2-VL-7B(Wang et al., 2024b) 33 144 15 54 246/1518 117 0 117/1518 72 825 897/1518 249 9 258/1518 1518/5472
MiniCPM-V 2.6(Yao et al., 2024) 150 14 16 30 210/1748 136 0 136/1748 61 1143 1204/1748 190 8 198/1748 1748/5624
@@ Appendix I Case Study
Here we show a terrible grounding problem issues during the inference of GPT4-o (Achiam et al., 2023). Due to the lengthy path, we only highlight the most critical subtask outputs and their execution results to emphasize errors (following cases are handled similarly). We can see that the ungrounded output directly prevents the process to be continued. Even after informing the model that the object doesn’t exist, the issue remains unresolved. The model continues to output incorrect objects or makes mistakes again after a few steps.
@@
Here we show a case of Qwen2-VL (Wang et al., 2024b) generating a large amount of meaningless outputs during the inference. As we can see, even though all subtasks are successful, the agent keep circling in place without making progress. These meaningless outputs quickly consume the remaining execution steps, ultimately causing the task to fail.
@@
Here we show how $D1$ error is solvable during the inference of Gemini-1.5-Pro (Team et al., 2024a). As we can see in step3, after a $D1$ error happens, a $Go\ to$ action can effectively solve it and facilitate the success of the whole trajectory.

### v2

@@ A.3 LLMs for Task Planning
Typical usages of LLM for task planing include treating LLM as a translator or a planner. There are also some studies combining LLMs with traditional PDDL [97, 98, 99, 100], in which LLM will be treated as a translator between the real-world and specific domain planner. But this method is limited by the performance of the domain planner and can’t leverage the powerful commonsense reasoning capabilities of LLMs to assist in planning. When LLM is treated as a planner, discrepancies between LLM’s outputs and real-world conditions always lead to execution failures. LLM-Planner [74] make a straightforward re-plan. Self-Refine [101] use single LLM as generator and evaluator. Reflexion [102] treats LLM as the Actor and the evaluator as the Critic. ViLA [103] utilizes GPT-4V [104] to obtain visual feedback. However, self-improvement relies heavily on prompt design and intrinsic capabilities of LLMs. If errors unrelated to planning occur, LLMs may struggle to self-correct. Inner Monologue [105] and RoCo [106] utilizes external collision detection and feedback. DoReMi [107] sets pre-defined constrains. Nevertheless, LLMs may make same mistakes in similar situations as the model weights are not changed. SayCan [73] trains a value function to consider both generated actions and their values. Remember [108] builds a memory module and retrieves similar state-action pairs. Retroformer [109] learns a retrospective model via policy gradient optimization. While RL-based adaptation mechanisms are able to adjust actions before execution, defining and training an effective value function or reward model is highly challenging. The recently popular DPO [24] algorithm greatly simplifies this process by requiring only a preference dataset. In our  HomieBot, we use DPO for model alignment, CoT [14] and self-reflection for decision-making. Additionally, error detection and feedback mechanisms are applied during low-level execution.
@@ B.1 Data Collection
We first randomly sample episode information provided by Replica Challenge [26] to build the task scenario, then we use the Fetch robot to collect EMMOE-100 in Habitat-lab v0.2.3. To facilitate data collection, we modify the original interaction script, and new interface can be seen in Fig. B1. The interface provides both third-person and first-person view observation to facilitate data collection, third-person observation is used to assist with collection, only first-person observation with 256*256 resolution will be saved. Notably, we only use the scene information to collect environment data, other functions and metrics in Replica Challenge are irrelevant to our work.
@@ E.3 Models
RT-1-X[67] RGB & Instructions Manipulation Picking & Placing
Octo[31] RGB & Instructions Manipulation Opening & Closing
NoMaD[32] RGB & Goal-Image Image-Navigation Navigate to Spot & Large Object
PixNav[33] RGB & Goal-Name Pixel-Navigation Navigate to Object
@@
[29] can flexible interact with target objects from various locations based on the integration of manipulative skills and mobility, while navigational skills are designed to accommodate multiple endpoints, ultimately leading to successful operations. Specifically, M3 implements these concepts by emphasizing mobile manipulation skills over fixed skills and training navigational skills using area targets rather than point targets.
@@
[30]) architecture utilizes image and text instructions as inputs, and generates discrete end-effector actions as outputs. Specifically, RT-1-X is a transformer-based model that guides robotic arms to complete various manipulation tasks. RT-1-X is an extension of the RT-1 ( [67]) model, which is designed for robot control and trained on a large-scale robot dataset.
@@
[31]) is an open-source, general-purpose policy for robotic manipulation based on transformers. It supports flexible task and observation definition and can be quickly integrated into new observation and action spaces.
@@
[32]) trains a single diffusion strategy for goal-oriented navigation and goal-independent exploration, the first one is to reach user-specified goals after localization and the second one is to search new environments. The method is instantiated using a transformer-based large-scale policy trained on data from various ground robots.
@@
[33]) is a pixel-guided navigational skill. It designs an LLM-based planner that utilizes common sense between objects and rooms to select the optimal waypoints, which are then executed by a pixel navigation strategy to achieve long-line-of-sight navigation. In this pipeline, we use its ability of finding the optimal waypoint and pixel navigation to navigate to some specific small object such as lemon and sugar box.
@@ F.1 SFT Augmentation
To expand the original dataset size, we first use GPT-4o [8] to regenerate text descriptions. Here is the regeneration code clip, we just show how to regenerate task descriptions, but the regeneration of subtask analysis uses the same template.
@@ G.1 Training Parameters
We use Video-LLaVA-7B [110] as our base model, we also use the training scripts they provide and partial parameters for $sft$ are as follows.
@@ H.1 Baseline Setup
Despite lowering the output format standards, the output from 7B-sized models still fails to meet our least requirements. They either do not output single-step subtasks or the subtask format is far from requirements. This issue is difficult to resolve by merely adjusting prompts. Therefore, we leverage the in-context learning abilities of these models by providing an output template example before each inference. Here, we provide the inference template for Qwen2-VL [39] MiniCPM-V 2.6 [40] respectively.
@@ H.2 Results
GPT-4o[8] 3.97 0.79 0.79 0 5.56 44.44 0 44.44 1.59 17.46 19.05 15.87 15.08 30.95 30.29
Gemini-1.5-Pro[37] 3.85 3.85 0 7.69 15.38 48.08 0 48.08 0 17.31 17.31 15.38 3.85 19.23 21.80
Qwen2-VL-7B[39] 0 0 0 0 0 100 0 100 0 0 0 0 0 0 20
MiniCPM-V 2.6[40] 0 0 0 0 0 100 0 100 0 0 0 0 0 0 6.67
o1[38] 0 0 1.18 10.06 11.24 21.89 0 21.89 0 48.52 48.52 18.34 0 18.34 20.56
@@
GPT-4o[8] 6.87 0.12 0.69 3.65 11.34 8.41 0.06 8.47 0.57 64.88 65.45 13.99 0.75 14.74 73.61
Gemini-1.5-Pro[37] 7.48 1.52 2.41 6.45 17.86 9.41 0 9.41 0 47.86 47.86 22.76 2.10 24.86 68.38
Qwen2-VL-7B[39] 2.17 9.49 0.99 3.56 16.21 7.71 0 7.71 4.74 54.35 59.09 16.40 0.59 17.00 27.74
MiniCPM-V 2.6[40] 8.58 0.80 0.92 1.72 12.01 7.78 0 7.78 3.49 65.39 68.88 10.87 0.46 11.33 31.08
o1[38] 1.16 0.07 1.96 4.86 8.05 8.64 0 8.64 0 56.75 56.75 26.27 0.29 26.59 51.77
@@
GPT-4o[8] 5 1 1 0 7/126 56 0 56/126 2 22 24/126 20 19 39/126 126/416
Gemini-1.5-Pro[37] 4 4 0 8 16/104 50 0 50/104 0 18 18/104 16 4 20/104 104/477
Qwen2-VL-7B[39] 0 0 0 0 0/9 9 0 9/9 0 0 0/9 0 0 0/9 9/45
MiniCPM-V 2.6[40] 0 0 0 0 0/1 1 0 0/1 0 0 0/1 0 0 0/1 1/15
o1[38] 0 0 2 17 19/169 37 0 37/169 0 82 82/169 31 0 31/169 169/822
@@
GPT-4o[8] 228 4 23 121 376/3317 279 2 281/3317 19 2152 2171/3317 464 25 489/3317 3317/4506
Gemini-1.5-Pro[37] 217 44 70 187 518/2900 273 0 273/2900 0 1388 1388/2900 660 61 721/2900 2900/4241
Qwen2-VL-7B[39] 33 144 15 54 246/1518 117 0 117/1518 72 825 897/1518 249 9 258/1518 1518/5472
MiniCPM-V 2.6[40] 150 14 16 30 210/1748 136 0 136/1748 61 1143 1204/1748 190 8 198/1748 1748/5624
@@
H.3 Task Performance
We also evaluate SR for each type of task defined in Section 2.2. As shown in Table H8, typical tasks are relatively easy due to straightforward processes and fewer overall steps. The most challenging are open-ended tasks, which usually have a very long total step count, with flexible processes and results, demanding powerful capabilities from both HLP and LLE models.
Table H8: Average success rate for each type of task. The highest value of each model is highlighted in bold.
Model Short- Horizon Long- Horizon Open- Ended Logical Human- Style
HomieBot (SFT) 43.75 24.60 18.52 34.01 25.24
HomieBot (SFT+DPO) 41.67 28.11 15.38 35.86 27.88
@@ Appendix I Case Study
Here we show a terrible grounding problem issues during the inference of GPT4-o [8]. Due to the lengthy path, we only highlight the most critical subtask outputs and their execution results to emphasize errors (following cases are handled similarly). We can see that the ungrounded output directly prevents the process to be continued. Even after informing the model that the object doesn’t exist, the issue remains unresolved. The model continues to output incorrect objects or makes mistakes again after a few steps.
@@
Here we show a case of Qwen2-VL [39] generating a large amount of meaningless outputs during the inference. As we can see, even though all subtasks are successful, the agent keep circling in place without making progress. These meaningless outputs quickly consume the remaining execution steps, ultimately causing the task to fail.
@@
Here we show how $D1$ error is solvable during the inference of Gemini-1.5-Pro [37]. As we can see in step3, after a $D1$ error happens, a $Go\ to$ action can effectively solve it and facilitate the success of the whole trajectory.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In Section 2, we propose EMMOE, the first unified benchmark designed to evaluate both high-level planners and low-level policies. In Section 2.2, we present the collection and features of EMMOE-100. In Section 2.3, we propose three novel metrics to complement existing evaluation methods. Next, we introduce our HomieBot and illustrate how its two main components HLP and LLE function in Section 3. In Section 4, we conduct experiments and demonstrate how to construct LMM-trainable SFT and DPO datasets and evaluate different levels of models. Finally, we conduct an in-depth analysis based on the detailed error information in Section 4.5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Section 5, we discuss following limitations and future works: Limited actions and available space in Habitat restrict the scope of task design. Besides, standardized output will sacrifice certain information precision. The growing number of model inferences will also lead to additional time costs. Moreover, we choose to conduct evaluations solely in simulation, where all researchers are required to make assessments under the same conditions, thus ensuring optimal fairness and consistency. In the future, we’ll collect more tasks, design a more efficient system, and explore real-world evaluations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We propose three new metrics in our work, and we have provide detailed definitions in Section 2.3, visible calculation processes in Appendix C, and complete evaluation codes in supplementary materials and project website.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: To ensure the reproducibility of our work, we provide data examples and visualizations in Appendix B and F, detailed training parameters and partial codes in Appendix H, entire codes in supplementary materials and project website. We also upload our models and datasets to the Huggingface platform.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have uploaded our codes on Github, our models and dataset on Huggingface, demonstrations on Youtube. we provide a project website at the end of the abstract, where all links can be found there.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, c, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 4.2, we provide some essential hyperparameter during model training, and full training settings in Appendix G. In Section 4.3, we introduce our metrics, baseline models, training and testing split selection.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification:
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Section 4.2 and Section 4.3, we discuss the type of compute workers GPU of model training and experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have carefully checked the codes we provide, to avoid potential harms and and better serve for the community.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Section 3, we also discuss the potential benefits of our system and setting. Though we evaluate on simulation, our methods have the potentials to facilitate to the real-world deployments and running. In Section 5, we discuss the limitations of our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our data is collected in an open-sourced simulator, and our models are also trained based on the open-source models. We have carefully follow the instructions of these simulators and models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have included citations of all works we used and mentioned and clearly mentioned the license and terms of use, which can be checked in the main paper and Appendix A.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have provided detailed and complete instructions and Readme file of our dataset, model, and codes in our supplementary materials and project website.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work doesn’t involve human participation, and doesn’t talk about the human-involved case either.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our work uses LLM for training and evaluations, as well a part of our proposed system. We have clearly pointed out what LLMs we use and their explicit functions in our work. We have also reported their performance and detailed analysis during experiments.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.